# Intramolecular Singlet Fission Through a Coherently Coupled Excimer-like Intermediate

### v1

@@ 2.4 Tracking electronic reorientation during $S_{1}\rightarrow[I]\rightarrow TT_{1}$
Normal mode analysis of the two most prominent modes seen on the $TT_{1}$ photoproduct suggest that multiple modes around $\sim$200-300 cm-1 correspond to inter-chromophore ruffling motions while those around $\sim$520-600 cm-1 correspond to inter-chromophore twisting motions (see Table S4), where both motions are expected to affect the intermolecular orbital overlaps. Joo and co-workers have shown that enhanced vibrational quantum beats on the excited state can directly report on the nuclear displacements that are the most displaced during internal conversion40, 41. Prominent inter-chromophore ruffling and twisting motions, observed as enhanced $\sim$210 cm-1 and $\sim$575 cm-1 quantum beats on the $TT_{1}$ photoproduct, suggests significant nuclear displacements along these coordinates during internal conversion, likely resulting in an excimer-like intermediate that then relaxes to a $TT_{1}$ state. Large admixtures of LECTLE-CT states created by orbital overlaps can lead35 to excimer-like states. Excimer formation48 in naphthalenediimide dimers is indeed sensitive to relative orientation and expected to be facile in case of eclipsed dimers. To further probe the nature of the intermediate $[I]$, we performed 2DES experiments and rate map analysis on the NDI dimer dissolved in a weakly polar solvent DCM solvent. The spectral features and observations are same as in the mixture of solvents (highly polar ACN with DCM). The corresponding 2DES spectrum at $T=$ 0.2 ps and the excitation wavelength dependence are shown in Figure S6 and suggest that the intermediate $[I]$ is likely to have only a weak $CT$ character not sufficient to significantly alter the wavelength dependent $TT_{1}$ formation kinetics. A functional role for modulation of inter-chromophore orbital overlaps was recently implicated49 in charge transport in NDI crystals. Low-frequency inter-chromophore motions and their effect on electron-phonon couplings are also known in pentacene polycrystalline films where Peierls couplings have been implicated16, 15, 50 in assisting SF in multiple reports. Our observations of enhanced excited state quantum beats along the inter-chromophore ruffling and twisting coordinates suggest a similar role of these nuclear motions in promoting iSF in the rigid NDI dimer studied here.
@@
The overall picture of iSF in NDI dimers that emerges from our observations is summarized in Figure 5c. The initial $\sim$200 fs rise of the intermediate followed by a slower rise of the ESA signal is reminiscent of the vibronically coherent $S_{1}$ + $TT_{1}$] intermediate formation reported in tetracene thin films26 by Friend and co-workers. Similar coherently coupled LECTLE-CT intermediates have been recently implicated25, 27 in symmetry-breaking charge transfer dyads. Interestingly, only minor changes in electronic anisotropy are seen during $T$ evolution of the $S_{1}$ state suggesting minimal electronic reorientation during the nuclear evolution that leads to the intermediate formation. The accompanying nuclear motions do not seem to localize the initially prepared state and maintain the strong electronic correlations between the sites (shared ground state). A complete localization of the initial photoexcitation would have resulted in GSB anisotropy increasing to $\sim$0.4 during intermediate formation. This suggests a picture where the $S_{1}$$TT_{1}$] intermediate is coherently coupled to the $LE$ state explaining its wavelength independent and fast formation – 30% of the $S_{1}$$TT_{1}$] intermediate is already formed within the IRF window in Figure 3d. Observations of upper 2D cross peak and reduced anisotropy in the GSB/SE band are also both consistent with the above picture. Notably, LETTLE-$T$ mixing with borrowed oscillator strength in the $TT_{1}$ state resulting in direct photoexcitation has been reported8, 7, 6, 5 in several acene thin films. Our observations show that similar effects may indeed be present in strongly coupled rigid iSF dimers even in the absence of thin films.
@@ Conclusions
The chemical space for iSF chromophores is rapidly expanding while detailed mechanistic insights into the electronic and nuclear motions involved during the correlated triplet pair formation are rare. NDI-based cyclophanes have been recently introduced28 as a new class of iSF systems where a combination of contortion and rotation makes iSF a favourable pathway. Through a combination of polarization-controlled 2DES and pump-probe spectroscopy, we have presented a comprehensive mechanistic investigation of the correlated triplet pair formation in these dimers with conclusive evidence for an excimer-like intermediate state whose $\sim$200 fs formation is excitation wavelength independent but subsequent $\sim$0.6-2 ps relaxation to $TT_{1}$ is strongly dependent on which Davydov component of the dimer is excited – a $\sim$3x faster triplet formation rate is observed when exciting the lower Davydov component. Enhanced vibrational quantum beats in the $TT_{1}$ photoproduct implicate inter-chromophore nuclear motions – twisting and ruffling – as key drivers of the intermediate formation and its subsequent relaxation to $TT_{1}$. Shared electronic correlations between the chromophore sites and surprisingly minimal electronic reorientation during these steps directly reported by 2DES cross-peaks and polarization anisotropy suggest a coherent picture of iSF where strong electronic correlations between the chromophores are maintained during the $S_{1}$ – $S_{1}$$T_{1}] – TT$$T_{1}$ electronic evolution. To best of our knowledge, the surprising lack of electronic reorientation during iSF has not been reported before and motivates refined electronic structure methods that account for this in predicting the high-spin state dynamics that follows $TT_{1}$ generation.
Taken together, this work represents the first experimental evidence of wavelength-dependent intramolecular singlet exciton proceeding through a coherently coupled excimer-like intermediate. Our findings advance the current understanding of iSF by highlighting the importance and providing a direct probe of, the shared electronic correlations between sites, enhanced nuclear motions during $TT_{1}$ formation and the electronic reorientation involved during $S_{1}\rightarrow[I]\rightarrow TT_{1}$ internal conversion. Our study also introduces broadband impulsive electronic polarization anisotropy51 as a powerful tool that can directly track electronic reorientation dynamics during singlet fission lending powerful mechanistic insights necessary for refining the current theoretical models for singlet fission.
Materials and Methods
2.5 Synthesis
All chemicals and solvents were purchased from Merck, Sigma-Aldrich, or TCI Chemicals (India) and used without further purification unless otherwise stated. Reactions for the NDI dimer synthesis were carried out in 4 mL heavy-walled Wheaton glass vials (Sigma-Aldrich).28 Thin-layer chromatography (TLC) was performed on aluminium-backed silica gel plates with a fluorescent indicator (Merck, Germany). The amino-substituted NDIs and the corresponding spectra of their radical ions have been reported earlier.29
Geometry optimization. Geometry optimization and frequency calculation of the NDI monomer and dimer structure was done using the B3LYP/6-31G level of theory using the Gaussian09 quantum chemistry package. From the analysis of the frequencies of the dimer it can be seen that there are several modes that are Raman active in the ground state. We have found that the modes around $\sim$ 200-300 cm-1 and $\sim$ 520-600 cm-1, as reported in Figures 5a,b involve complete skeletal movements as denoted in Figure 5c.
Optical Setup
The experimental setup, data collection and data analysis details have been published elsewhere52, 53 and described in detail in Section S1. Briefly, the 2DES as well as the PP setups are based on a YAG generated white light continua (WLC) separately generated in the pump and probe arms. Using a combination of 100 kHz shot-to-shot detection with rapid stage scan we have previously reported the spectrometer sensitivity and throughput to be comparable or better than the equivalent acousto-optic pulse shaping based approaches. Optical dispersion is minimized by a combination of chirped mirrors and glass wedges for maximizing the two-photon intensity autocorrelation signal. The resulting FWHM duration of the instrument response function, obtained after a Gaussian instrument response convoluted global fit54 across the $\sim$200 nm probe spectral bandwidth, is 38 fs with a probe rise time dispersion that is within the signal-to-noise ratio (SNR) of the measurement (Figure S1a). Signal fluctuations due to this sample movement are minimized by optimizing the probe spectrum fluctuation at the CCD detector at $T$ where pump-probe signal is maximum.
For polarization anisotropy experiments, the probe vertical polarization (V) extinction ratio at the sample location is $\sim$1430. The extinction ratio for the pump polarization at the sample location was determined to be $\sim$4498. Together, these settings reproduce the early $T$ electronic anisotropy of 0.396 $\pm$ 0.018 for oxazine 720 in methanol ( Figure S9), as expected for an isolated transition dipole. Due to coherent vibrational wavepackets, the anisotropy is averaged over a 200 fs $T$ range around 500 fs. Note that the polarization dependent reflection of the focusing parabolic mirror was also considered in the above calculation. However, possible polarization dependent changes in reflected bandwidth are ignored.
[…]
The linear excitation versus emission maps are collected using Horiba Fluoromax at 700 nM sample concentration. No fluorescence re-absorption effects are expected at this low concentration because the average number density is much smaller than the Förster critical radius for energy transfer. The absence of reabsorption effects is further confirmed in Figure S2b.
Data Availability
All data required for interpretation and verification of experimental results and reproducing the theoretical model are given in the paper and the Supplementary Information.
Code Availability
Codes for the exciton models and the resulting polarization anisotropy presented in the paper are available from the corresponding author upon reasonable request.
Author Contributions
V.T. conceptualized the project. C.M.H., V.P.S. and P.M. synthesized the samples. S.P. and A.B. performed all the 2DES and pump-probe experiments. S.P. performed all the data analysis and theoretical calculations of polarization anisotropy. S.S. and D.G. performed geometry optimization and calculations of Raman activity of the dimer. V.T. and S.P. wrote the manuscript with inputs from all the authors.
@@ Acknowledgements
A.B. acknowledges the Prime Ministers’ Research Fellowship, MoE, India. S.P. acknowledges the MHRD research fellowship from the Indian Institute of Science. P.M. acknowledges SERB grant no. CRG/2021/008494 for partial funding. P. M. thanks Dr. Deepak Bansal for his valuable initial inputs in the synthesis of the NDI dimers. D.G. acknowledges the support of the SERB-POWER Fellowship (SPF/2021/000194) and the computational facilities provided by IACS. S.S. acknowledges funding support from the DST-INSPIRE program through a Junior Research Fellowship. V.T. acknowledges funding from ANRF under grant sanction number CRG/2023/000327, from ISRO-STC under grant sanction number ISTC/CSS/VT/505, and from the Quantum Research Park (QuRP) funded by the Government of Karnataka.
Competing Interests
The authors declare no competing interests.

### v3

@@ 2.4 Tracking electronic reorientation during $S_{1}\rightarrow[S_{1}+TT_{1}]\rightarrow TT_{1}$
Normal mode analysis of the two most prominent modes seen on the $TT_{1}$ photoproduct suggest that multiple modes around $\sim$200-300 cm-1 correspond to inter-chromophore ruffling motions while those around $\sim$520-600 cm-1 correspond to inter-chromophore twisting motions (see Table S4). Both motions are expected to affect the intermolecular orbital overlaps. Joo and co-workers have shown44, 45 that enhanced vibrational quantum beats on the excited state can directly report on the nuclear displacements that are the most displaced during internal conversion. Prominent inter-chromophore ruffling and twisting motions, observed as enhanced $\sim$210 cm-1 and $\sim$575 cm-1 quantum beats on the $TT_{1}$ photoproduct, suggests significant nuclear displacements along these coordinates during internal conversion, likely resulting in an excimer-like $S_{1}+TT_{1}]$ intermediate which then relaxes to the $TT_{1}$ state. Excimer formation53 in naphthalenediimide dimers is indeed sensitive to relative orientation and expected to be facile in case of eclipsed dimers. A functional role for modulation of inter-chromophore orbital overlaps was recently implicated54 in charge transport in NDI crystals. Low-frequency inter-chromophore motions and their effect on electron-phonon couplings are also known in pentacene polycrystalline films where Peierls couplings have been implicated16, 15, 55 in assisting SF in multiple reports. Notably, similar low frequency ($\sim$ 200 cm-1) intermolecular out-of-plane vibrations have also been implicated23, 24 in excimer intermediate formation in perylenebisimide (PBI) dimers. Similarly, $\sim$500cm-1) ring breathing and deformation motions have been implicated21 in the iSF process in PBI trimers. Our observations of enhanced excited state quantum beats along the inter-chromophore ruffling and twisting coordinates suggest that such structural dynamics may be more general across various iSF platforms.
Large admixtures of LECTLE-CT states created by orbital overlaps can lead39 to excimer-like states. To further probe the $CT$ character in the intermediate, we performed 2DES experiments and rate map analysis on the NDI dimer dissolved in a non-polar (weakly polar) DCM solvent. The resulting spectral features and observations are same as in the polar solvent mixture (highly polar ACN with DCM). The corresponding 2DES spectrum at $T=$ 0.2 ps and its excitation wavelength dependence are shown in Figure S6, and suggest that the $S_{1}+TT_{1}]$ intermediate is likely to have only a weak $CT$ character not sufficient to significantly alter the wavelength dependent $TT_{1}$ formation kinetics. Excimer intermediate in SF with weak to absent $CT$ character has been previously reported26, 46. However, excimer-mediated SF has also been reported in systems where the $CT$ character in the excimer intermediate is tuned through molecular packing27 and solvent polarity21 and larger $CT$ character in the excimer intermediate is correlated with faster $TT_{1}$ formation. Solvent polarity independent rates suggest only weak $CT$ character is involved and therefore we assign the intermediate to an excimer-like species where we confirm the excimer-like nature of the intermediate directly by analyzing the nuclear motions corresponding to the quantum beats that are enhanced in the $TT_{1}$ photoproduct, an approach suggested44, 45 by Joo and co-workers.
@@
The overall picture of iSF in NDI dimers that emerges from our observations is summarized in Figure 5c. The initial $\sim$200 fs rise of the intermediate followed by a slower rise of the ESA signal is reminiscent of the vibronically coherent $S_{1}$ + $TT_{1}$] intermediate formation reported in tetracene thin films30 by Friend and co-workers. Similar coherently coupled LECTLE-CT intermediates have been recently implicated29, 31 in symmetry-breaking charge transfer dyads. Interestingly, only minor changes in electronic anisotropy are seen during the evolution of the $S_{1}$ state suggesting minimal electronic reorientation during the nuclear evolution that leads to the intermediate formation. The accompanying nuclear motions do not seem to localize the initially prepared state and maintain the strong electronic correlations between the sites (shared ground state). A complete localization of the initial photoexcitation would have resulted in GSB anisotropy increasing to $\sim$0.4 during intermediate formation. This suggests a picture where the $S_{1}$$TT_{1}$] intermediate is coherently coupled to the $LE$ and $TT_{1}$ states with strong correlations between the sites, explaining its wavelength independent and fast formation – 30% of the $S_{1}$$TT_{1}$] intermediate is already formed within the IRF window in Figure 3d. Observations of upper 2D cross peak and reduced anisotropy in the GSB/SE band are also both consistent with the above picture. Notably, LETTLE-$T$ mixing with borrowed oscillator strength in the $TT_{1}$ state resulting in direct photoexcitation has been reported8, 7, 6, 5 in several acene thin films. Our observations show that similar effects may indeed be present in strongly coupled rigid iSF dimers even in the absence of thin films. Tracking electronic dynamics accompanying structural evolution directly shows that the $TT^{1}$ photoproduct maintains significant singlet character which likely accelerates its fast recombination.
@@ CONCLUSIONS
The chemical space for iSF chromophores is rapidly expanding while detailed mechanistic insights into the electronic and nuclear motions involved during the correlated triplet pair formation are rare. NDI-based cyclophanes have been recently introduced32 as a new class of iSF systems where a combination of contortion and rotation makes iSF a favourable pathway. Through a combination of polarization-controlled 2DES and pump-probe spectroscopy, we have presented a comprehensive mechanistic investigation of the correlated triplet pair formation in these dimers with conclusive evidence for an excimer-like intermediate state whose $\sim$200 fs formation is excitation wavelength independent but subsequent $\sim$0.6-2 ps relaxation to $TT_{1}$ is strongly dependent on which Davydov component of the dimer is excited – a $\sim$3x faster triplet formation rate is observed when exciting the lower Davydov component. Enhanced vibrational quantum beats in the $TT_{1}$ photoproduct implicate inter-chromophore nuclear motions – twisting and ruffling – as key drivers of the intermediate formation and its subsequent relaxation to $TT_{1}$. Shared electronic correlations between the chromophore sites and surprisingly minimal electronic reorientation during these steps directly reported by 2DES cross-peaks and polarization anisotropy suggest a coherent picture of iSF where strong electronic correlations between the chromophores and singlet-triplet mixing is maintained during the $S_{1}$ – $S_{1}$$T_{1}] – TT$$T_{1}$ electronic evolution. The surprising lack of electronic reorientation during iSF in strongly coupled dimers has not been reported before and motivates synthetic design and theoretical models which balance strong correlations between the triplets while diminishing singlet-triplet mixing with nuclear evolution. Predictions of high-spin state dynamics arising from correlated triplets also need to account for singlet-triplet mixing during $TT_{1}$ relaxation. In the context of tracking21 structural dynamics during iSF, our observations demonstrate that tracking electronic dynamics provides complementary synthetic and theoretical insights for advancing iSF systems. Our study also introduces impulsive electronic polarization anisotropy56 as a powerful tool that can directly track electronic reorientation dynamics during singlet fission lending powerful mechanistic insights necessary for refining the current theoretical models for singlet fission.
ASSOCIATED CONTENT
Supporting Information
Experimental details, including synthesis and geometry optimizations, optical layout, spectroscopic measurement procedures, as well as wavelength-dependence analysis.
Notes The authors declare no competing interests.
@@ ACKNOWLEDGEMENTS
V.T. thanks Prof. Jyotishman Dasgupta, Department of Chemical Sciences, TIFR Mumbai for initiating this collaboration and for the numerous fruitful discussions. A.B. acknowledges the Prime Ministers’ Research Fellowship, MoE, India. S.P. acknowledges the MHRD research fellowship from the Indian Institute of Science. P.M. acknowledges SERB grant no. CRG/2021/008494 for partial funding. P. M. thanks Dr. Deepak Bansal for his valuable initial inputs in the synthesis of the NDI dimers. D.G. acknowledges the support of the SERB-ANRF Fellowship (CRG/2023/001806) and the computational facilities provided by IACS. S.S. acknowledges funding support from the DST-INSPIRE program through a Junior Research Fellowship. V.T. acknowledges funding from SERB-ANRF under grant sanction number CRG/2023/000327, from ISRO-STC under grant sanction number ISTC/CSS/VT/505, and from the Quantum Research Park (QuRP) funded by the Government of Karnataka.